Zephyr Farming Pty Ltd is considering the purchase of a wind turbine generator in order to generate electricity and to reduce the electricity costs for their offices, which are located in Toowoomba. Currently the business uses 60,000 kilowatt hours (kWh) per quarter (3 months) at an average cost of $0.30 per kwh, supplied by the local coal fired power station. The current required rate of return used to evaluate projects is 6%, with a required payback period of 3 years.
The Queensland government started a scheme to provide an incentive for business to use alternative sources of renewable power. The incentives are 5% immediate reimbursement of the purchase and installation costs. This reimbursement can be paid to the supplier providing and installing the equipment. Therefore, the net cash flow from Zephyr Farming Pty Ltd is the purchase cost plus the installation cost less the 5% incentive back from government to offset these costs.
Project details:
Cost of wind turbine generator          $6,000
Cost to install turbine and generator (by supplier)        $450
Expected cash incentive back from government to offset cost of the panels paid immediately the wind turbine generator installed           5% of total costs
Turbine expected (on average) generated kilowatt              300 hours per month
Generator’s expected life (in years)                   15 years

Mainly answer the question 6,7 and 8 please. Thank you.

Requirements
1. Calculate the total initial investment and the net annual savings from installing the wind generator. (8 marks)
2. Calculate the payback period of the wind generator. (4 marks)
3. Comment on whether this project meets the company’s Payback Period requirements and whether you should base your decision only in the payback period calculation.
(1 mark)
4. Calculate the net present value (NPV) of the project with a required rate of return at 6%. As the savings are the same each year, use the present value table on the next page for an ordinary annuity. (9 marks)
5. Comment on whether Zephyr should proceed with this project using the Net Present Value method and whether you should base your decision only in this quantitative calculation. (1 mark)
6. Comment on two limitations with the payback period method after considering your answers to 2 and 4 above (2 marks)
7. What qualitative, non-financial factors (not included in the quantitative analysis) would influence your opinion on whether the investment in the wind generator should proceed. (2 marks)
8. As the Chief Operating Officer (COO) of Zephyr Farming Pty Ltd, provide [in an internal informal report email] your recommendation(s) about what to implement [use your answers in 1 to 7 to support your recommendation(s)]. You report will be sent directly to Hoang Binh, who is the Chief Financial Officer (CFO). (3 marks)
Answer #1

1. Cost of turbine generator = $6,000

Add: installation cost = $450

Less: Reimbursement from government = 5%*6,450 = $322.5

Total initial investment = $6,127.5

Calculation of net annual savings:

Generation of kilowatt hours per month = 300

Annual = 300*12 = 3,600

Cost per kWh = $0.30

Net annual savings = 3,600*0.30 = $1,080

2. Payback period is the time period in which the initial cost is recovered.

Payback period of generator = 6,127.5/1,080

= 5.67 years

3. The acceptable payback period for company is 3 years, while the payback period for generator is 5.67 years. Hence, the generator doesn't meet company's payback period requirements.

It is not right to base the decision on only payback period. The generator will provide benefits for 15 years and required return is 6% on investment.

Based on this information, the generator will provide benefits for a long term and hence is a good investment.

4. NPV = Present Value of Cash inflows - Present Value of Cash Outflows

= 1,080*PVAF(6%, 15 years) - 6,127.5

= 1,080*9.712 - 6,127.5

= 4,361.46

5.Since NPV of the project is positive, Zephyr should proceed with this project as per NPV method.

Payback period of investment should also be considered. Also, factors such as availability of funds etc. should also be considered.

6.Limitations of payback period:

1.It does not consider inflows generated after payback period.

2. It does not take into account time value of money

7.Qualitative non-financial factors are:

a.services provided by vendor

b.Warranty on purchase

c.Capability of employees to operate the generator.

8.Report:

The payback period of generator is higher than the acceptable payback period. However, the investment will generate a positive NPV and cost savings for 15 years. Investment seems goods from financial point of view. The generator should be purchased after taking into account the qualitative factors such as feasibility of operation, service provided by the vendor etc.